OpenStudy (anonymous):

second order partial derivative of 2e - 2e^2y

OpenStudy (anonymous):

y' = 2e - 2e^2y + 4e^2y =2e + 2e^2y y" = 2e - 2e^2y + 4e^2y + 4e^2y + 8e^2y = 2e + 14e^2y

OpenStudy (unklerhaukus):

/??

OpenStudy (anonymous):

@gerryliyana r u taking the derivative w.r.t. y??????/

OpenStudy (anonymous):

nope.., thats wrong..,

OpenStudy (anonymous):

according to me, if we are taking derivative w.r.t. y then y'= 0-2(e^2y)*2 = -4e^2y y''= -8e^2y

OpenStudy (anonymous):

yes y' = -4e^2y ln(e) y" = -8*e^(2*y)*ln(e)^2 right??

OpenStudy (anonymous):

@gerryliyana what does ln(e) signify in y' and ln(e)^2 signify in y''?????

OpenStudy (anonymous):

hei @niksva relax..,hehe y' = -4e^2y ln(e) = -4e^2y (1) = -4e^2y y" = -8*e^(2*y)*ln(e)^2 =-8*e^(2*y) (1^2) =-8*e^(2*y) right?? hint. ln(e) =1

OpenStudy (anonymous):

Ok., @Rapi This is the complete solution: \[u = 2e -2e ^{2y}\] for first derivative: \[\frac{ d }{ dy} (2e -2e ^{2y})\] differentiate the sum term by term and facetor out constant \[\frac{ du }{ dy } =2\left( \frac{ d }{ dy } (e) \right) - 2e ^{2}\left( \frac{ d }{ dy } (e^{2y}) \right)\] derivative of e is zero \[\frac{ du }{ dy } =2(0) - 2e ^{2}\left( \frac{ d }{ dy } (e^{2y}) \right)\] using the chain rule, where s =2y \[\frac{ d }{ dy } (e^{2y}) = e ^{s} \frac{ ds }{ dy }\] and \[\frac{ de ^{u} }{ u}=e ^{u}\] \[-2e ^{2y}\left( 2\left( \frac{ d }{ dy } \left( y \right) \right) \right)\] factor out contans, so u You will get the answer for second derivative \[\frac{ d ^{2}u }{ dy ^{2} } =\]

OpenStudy (anonymous):

for second derivative \[\frac{ d }{ dy } (-4e ^{2y}) \] factor out constants \[-4\left( \frac{ d }{ dy } (e ^{2y}) \right)\] using the chain rule \[\frac{ d }{ dy } (e ^{2y}) = e ^{S}\frac{ dS }{ dy } \] where S =2y and \[\frac{ de ^{S} }{ dS} = e ^{S}\] becomes \[-4\left( e ^{2y} \left( \frac{ d }{ dy }(2y) \right)\right)\] Factor out constants \[-4e ^{2y}\left( 2\left( \frac{ d }{ dy }(y) \right) \right)\] How is the result of a derivative of y?? Good luck @Rapi!!